OpenStudy (superhelp101):

Question:

OpenStudy (superhelp101):

I really don't understand what the symbol between the P and Q is

OpenStudy (anonymous):

its multiplication

OpenStudy (superhelp101):

yeah, how do i do it

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

they invented a new operation

OpenStudy (anonymous):

p ? q = (p*q) / ( p - q )

OpenStudy (superhelp101):

okay what does P and Q equal

OpenStudy (anonymous):

start with 6 and 4

OpenStudy (anonymous):

We can do it in steps. First simplify the terms in the parentheses ( 6 ? 4 ) = ( 6 * 4 ) / ( 6-4)

OpenStudy (superhelp101):

okay i did it

OpenStudy (anonymous):

( 6 ? 4 ) = ( your answer ) ? 3

OpenStudy (superhelp101):

12 ? 3

OpenStudy (anonymous):

yes

OpenStudy (superhelp101):

yes we would divide next to get 4. I don't know whythou

OpenStudy (anonymous):

( 12 * 3 ) / (12 - 3) = 36/9 = 4

OpenStudy (superhelp101):

omg I get know thank you so much

OpenStudy (anonymous):

Your welcome